Pls help solve the equation by factoring, kindly show the complete solution. Thanks much.
x^2-2x+2=0

Answer by jorel1380(3719)

x²-2x+2=0
x²-2x+1+2=1
x²-2x+1=-1
(x-1)²=-1
x-1=i
x=i+1
☺☺☺☺